OpenStudy (anonymous):

Which point lies in the solution set for the following system of inequalities? x − y > 3 x + 2y < 6 (7, 1) (3, 4) (−2, 2) <- my answer plz correct me if im wrong (4, −2)

hartnn (hartnn):

not correct.

hartnn (hartnn):

start plugging in points in the inequalities say 3,4 3-4>3 -1 >3 FALSE so, 3,4 is also not correct

hartnn (hartnn):

can you do that same thing for 7,1 ?

hartnn (hartnn):

and 4,-2 ?

OpenStudy (anonymous):

i will try im not so good at this lesson so 4,-2 4-(-2)>4 did i do it right

OpenStudy (anonymous):

sorry im not good at algebra

hartnn (hartnn):

correct, go on!

hartnn (hartnn):

4- (-2) = 4+2 = ...

OpenStudy (anonymous):

what do I do next ?

hartnn (hartnn):

6>4 thats true, right ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

also check 2nd equation, x+2y <6 4 +(2(-2)) <6 4-4 < 6 0<6 this is also true, right ?

OpenStudy (anonymous):

Thank you so much :)

hartnn (hartnn):

since both are true, 4,-2 is your answer! :)

hartnn (hartnn):

welcome ^_^